Eagle Scout Leadership Service Project Workbook

David Fifield

Scout’s name

16816 E. Kenyon Dr. Aurora, CO 80013

Address

(303) 766–3525

Telephone No.

Troop 171

Unit No.

Centennial

District

Denver Area

Local council

John Gonsalves

Vivian Coleman PROJECT DESCRIPTION

Unit leader’s name Address Telephone No.

Unit advancement committee person’s name Address Telephone No.

Describe the project you plan to do.

I plan to collect books for a local library. I will accept both donations of books and money towards the purchase of new books. I hope to promote this project through the distribution of flyers, announcements made at places like Trinity Baptist Church, advertisements in local publications, and the word-of-mouth of interested Scouts, whose parents and friends can make donations. Additionally, I plan to build a web site for this project, which will contain all the project’s important information. Through the web site, people will be able to access the most up-to-date information and will have a convenient way to contact me with questions or donations.

One of the important goals of this project is to collect books related to computers and technology, these being personal interests of mine. I hope to make the benefitting library’s technology section both current and complete. Because of the changing nature of technology, many of these books will have to be purchased new. This will be done with money collected from donations and from fundraisers.

If the library wishes, Boy Scouts could be recruited to label, sort, and repair the books collected. What group will benefit from the project?

Aurora Public Library (South Branch)  (303) 693–1440 Name of religious institution, school, or community  Telephone No.
15324 E. Hampden Cir.  Aurora  CO  80013 Street address  City  State  Zip code

My project will be of benefit to the group because:

The library is something that benefits the community. Therefore, anything that benefits the library benefits the community. I was inspired to do this project after visiting the library to find some information, and being unable to find what I sought. I think that a more complete library would be better able to serve the needs of people seeking knowledge. This concept was discussed with my unit leader on \frac{10 Apr 2000}{Date}

The project concept was discussed with the following representative of the group that will benefit from the project.

Barbara Schneller  05 Apr 2000
Program Librarian  (303) 693–1440

Representative’s name  Date of meeting Representative’s title  Phone No.

PROJECT DETAILS

Plan your work by describing the present condition, the method, materials to be used, project helpers, and a time schedule for carrying out the project. Describe any safety hazards you might face, and explain how you will ensure the safety of those carrying out the project. If appropriate, include photographs of the area before you begin your project. Providing before-and-after photographs of your project area can give a clear example of your effort. The Present Condition:

To describe the present condition of the library’s technology section, I simply counted books using the library’s indexing system. I can evaluate my project’s success simply by comparing this number to the number of books my project provides.

There are approximately three shelves of books of the type I want to collect. These are found in sections 004–006. According to the library’s computer records:

Number of books in section 004: 354
Number of books in section 005: 450
Number of books in section 006:  48
Total: 852

Of these books, 655 were published since 1994. Also, there are about 30 books on computer repair, in section 621. The Method:

First of all, I will accept books and monetary donations personally, through flyers and the project’s web site. Information regarding the project will be distributed by me, interested Scouts, and announcements at Queen of Peace Catholic Church and the pancake breakfast fundraiser, described below.

In order to demonstrate my leadership ability, I will coordinate and hold a pancake breakfast, with the assistance of Troop 171. The proceeds from this fundraiser will be used to purchase new books for the library.

In total, I hope to collect at least 250 books, at least 50 of which relate to technology and are sufficiently recent, and $200 in donations. Materials to be Used:

To collect donations and transport materials, I will use my personal vehicle.

For the fundraiser, Troop 171 has offered to donate the stoves and the facility will most likely be Trinity Baptist Church. Disposable plates, silverware, etc., will be donated by my family. Project Helpers:

I will be assisted by Boy Scout volunteers from Troop 171. They will provide most of the labor for the pancake breakfast and will be responsible for much of the flyers’ distribution.

Adults from the Troop will cook the food. Time Schedule:

01–02 Jul 2000 — Announcement & flyers at Queen of Peace church
26 Aug 2000 — Pancake breakfast
02 Sep 2000 — Final donation of books to the library

“Before” Photographs (None)

Approval Signatures for Project Plan Project plans were reviewed and approved by:

Religious institution, school, or community representative  Date
Scoutmaster/Coach/Advisor  Date
Unit committee member  Date
Council or district advancement committe member  Date

IMPORTANT NOTE: You may proceed with your leadership project only when you have… — Completed all the above mentioned planning details

— Shared the project plans with the appropriate persons

— Obtained approval from the appropriate persons CARRYING OUT THE PROJECT

Record the progress of your project. Keep a record of how much time you spend planning and carrying out the project. List who besides yourself worked on the project, the days they worked, the number of hours they worked each day, and the total length of time others assisted on the project. If appropriate, list the type and cost of any materials required to complete the project. If your original project plan changes at any time, be sure and document what the change was and the reason for the change. Hours I Spent Working on the Project The length of time spent should be as adequate as is necessary for you to demonstrate your leadership of two or more individuals in planning and carrying out your project. Project Log

22 Aug 2000 1:00  Wrote letter to present to corporate donators
22 Aug 2000 1:15  Went seeking corporate donations
25 Aug 2000 0:15  Bought pancake mix and orange juice
26 Aug 2000 1:00  Created flyer for pancake breakfast
28 Aug 2000 0:30  Made copies of the pancake breakast flyer
28 Aug 2000 0:15  Received gift certificate from Cub Foods
29 Aug 2000 0:40  Got materials from trailer
31 Aug 2000 0:25  Got the eggs
01 Sep 2000 2:00  Bought supplies for project
02 Sep 2000 6:00  Eagle project
08 Sep 2000 0:40  Brought books to library

Hours I spent… Planning the project: 7:20 Carrying out the project: 6:40 Total hours I spent working on the project: 14:00 Hours Spent by Scouts, Venturers, or Other Individuals Working on the Project

Name  Date  No. of Hours
Mary Kay Brengosz  01 Sep 2000  2:00
David C. Fifield  02 Sep 2000  5:15
Mary Kay Brengosz  02 Sep 2000  5:15
Shelly Czarnek  02 Sep 2000  5:00
Thomas Lambrick  02 Sep 2000  3:20
Andy Horn  02 Sep 2000  4:11
Matt Ingleby  02 Sep 2000  3:00
Tom Ingleby  02 Sep 2000  5:00
Kurt Horn  02 Sep 2000  4:16
Marilu Hirschinger  02 Sep 2000  5:15
Molly Doubrava  02 Sep 2000  2:00
Justin Doubrava  02 Sep 2000  2:15
Curtis Fendel  02 Sep 2000  4:45
Zach Czarnek  02 Sep 2000  5:00
Josh Munroe  02 Sep 2000  5:15
Jason Doubrava  02 Sep 2000  2:00
Mark Perez  02 Sep 2000  3:15
Casey Jenkins  02 Sep 2000  3:15
Scott Jenkins  02 Sep 2000  3:15
Linda Jenkins  02 Sep 2000  3:15
Al Jenkins  02 Sep 2000  3:15
Joe Cutshall  02 Sep 2000  1:30
Jean Cutshall  02 Sep 2000  1:30
Matt Hirschinger  02 Sep 2000  2:15
Mark Coleman  02 Sep 2000  2:05
Matt Deall  02 Sep 2000  2:00
David Dunn  02 Sep 2000  1:41
Thomas Deall, Jr.  02 Sep 2000  2:00
Joseph Deall  02 Sep 2000  2:00
Samantha Deall  02 Sep 2000  2:00
Lisa Deall  02 Sep 2000  2:00
Tom Deall  02 Sep 2000  2:00
Zack Hirschinger  02 Sep 2000  0:35
Penny Ruddell  02 Sep 2000  1:15
Bryan Ruddell  02 Sep 2000  1:15
Tom Mentzer  02 Sep 2000  3:45

Total number of hours others worked on the project: 107:53 For a grand total, add the total number of hours you spent on the project to the total number of hours others worked on the project: 121:53 Materials Required to Complete the Project

Type of Material  Cost of Material
Juice and pancake mix  $24.83
Coffee pot rental  $18.00
Pancake syrup  $4.99
Plates  $4.39
Coffee  $4.10
Silverware  $2.50
Creamer  $2.49
Napkins  $1.45
Cups  $2.18
Sugar  $0.80
Butter  $0.60
Eggs  $0.00
Profits/Donations  Amount
King Soopers gift certificates  $20.00
Cub Foods gift certificate  $10.00
Eagle project donations  $260.10
Other donations  $50.00

My total costs were $66.33. My total income was $340.10, creating a profit of $273.77. Changes List any changes made to the original project plan and explain why those changes were made.

There were no announcements made in publications, for a lack of funds and time. For the Spanish mass at Queen of Peace, I produced a Spanish translation of the main flyer. The library preferred to sort and catalog the books for itself, so no Scouts participated in this.

The pancake breakfast was held September 2nd. The books were donated the following Friday, September 8th, to work with the library’s schedule. Summary

I consider my project a success. I exceeded my goals with only the few changes in plan noted above. If I had this project to do over, however, I would have done a few things differently. I would have advertised the breakfast better. Far fewer people showed up than I expected, but they were more generous than I expected, too. I would have sought more corporate donations. While I was able to keep costs fairly low, I might have been able to cover all my expenses without using donations had I visited a few more stores.

I am especially grateful to the members of Troop 171 who provided so much assistance, my Eagle advisors David Ernst and Chris Easton, and the staff of the Aurora Public Library, especially Program Librarian Barbara Schneller.

Photographs “After” Photographs Including photographs of your completed project (along with the “before” photographs with the project description) helps present a clearer overall understanding of your effort.

These pictures and more, along with a comprehensive summary of the project, may be

found on the web at 171 Recon’s web site: http://171recon.freeservers.com/outings/.

Approvals for Completed Project Start date of project: 10 Aug 2000 Completion date of project: 08 Sep 2000 The project was started and has been completed since I received the Life Scout rank, and is respectfully submitted for consideration.

Applicant’s signature  Date

This project was planned, developed, and carried out by the candidate.

Signature of Scoutmaster/Coach/Advisor  Date
Signature of the representative of religious institution, school, or community  Date